What is the probability of landing on number 4 if you spun a spinner numbered 1 through 8?
1/8 or .125 or 12.5%
Moment generating function of a negative binomial distribution?
[(1 - p)/(1 - pet)]r for t < -ln(p)
where
p = probability of success in each trial,
r = number of failures before success.
3/16 * 7/15 = 21/240
10/3
1/2
What is the probability of choosing a face card on the second draw if the first draw was a king?
7/26
Why were people attracted to Judaism?
Everyone is attracted to different philosophies for different reasons. It really can't be summed up in a generic statement.
In the early days of monotheism, people might have been attracted to it because it was new and different.
In the Rabbinic age (arguably up until modern times), outsiders were historically NOT attracted to Judaism because the Jewish people were persecuted and hated.
What is the probability of a die landing with a 3 up?
1 in 6, with a one-sided die. You have 6 possible outcomes, all have an even chance of happening. Therefore rolling a three is a 1 in 6, or a 1/6 chance.
Which shows the probability of choosing 2 blue socks from a drawer having 7 blue and 3 white socks?
The probability of randomly choosing 1 blue sock is 7/10. The probability of randomly choosing 2 blue socks in a row is 7/10 x 7/10 = 49/100.
Which sample is more representative of the population?
We could answer that in a snap, if we knew something about the population
and could see the list of proposed samples. Call me crazy, but I'm suspecting
wherever you copied the question from, all that stuff was right there with it.
Are heteroscedasticity robust standard errors consistent?
The answer depends on consistent with WHAT!
What is the probability you spin a number that is is not a perfect square?
depends on the numbers on the spinner. if 1 thru 10, 7/10
How do you calculate geometric probability?
In the simplest case, a geometric probability is one that is given in terms of the ratio of two areas. For example, suppose a parachutist could land anywhere on a 10 square kilometre area of open country with equal probability, and you wanted to know how probable it would be that the parachutist would land on a designated area of 2 square kilometres with that part of open country.
Then the probability would be 2 / 10 = 0.2
The same principles apply in more and more difficult or complex cases, and in spaces of higher dimension. For instance, one can discuss geometric probabilities involving three-dimensional space.
11 x 10 x 9 x 8 x 7 x 6 = 332,640.
How common is it to have identical twins?
Well, as an identical twin myself, I believe there is a one out of twelve chance of having identical twins, although, I'm not 100% sure....
If something has probability 1000 percent is it sure to happen?
In the context of the usage 1000%, it would be sure to happen. As far as probability is concerned, the probability of certain is 1 or 100%.
Is it possible to have 100 percent confidence interval?
Yes. Simply make sure that the interval is greater than or equal to the range of the random variable.
What are the odds against getting exactly two heads in three successive flips of a coin?
Three in eight are the odds of getting exactly two heads in three coin flips. There are eight ways the three flips can end up, and you can get two heads and a tail, a head and a tail and a head, or a tail and two heads to get exactly two heads.
What is the probability of rolling an eight with one roll of a single number cube?
You can not roll an 8 on a 6 sided cube. The chances would be 0%, 0/0, or 0 to 0
Using the Poisson approximation, the probability is 0.0418
What is the probability that the top card of a deck is a heart or a 2?
With a well shuffled normal deck, it is 16/52 or 4/13.